Angelena files as a head of household. In 2019, she reported $54,250 of taxable income, including a $12,400 qualified dividend. What is her gross tax liability, rounded to the nearest whole dollar amount? (Use the tax rate schedules, long-term capital gains tax brackets)
Answer #1

Answer:

Yax Year = 2019

Filing status = Head of Household

Taxable income = $54,250

Ths includes qualified dividend = $12,400

Tax on ordinary income of (54250 - 12400 =) $41,850

= 13850*10% + (41850 - 13850) * 12%

= $4745

For Head of Household, long term capital gain tax rate for bracket $0 to $52,750 = 0%

Tax on qualified dividend = 12400 * 0% = $0

Her gross tax liability = 4745 + 0 = $4745

Her gross tax liability = $4,745